1. Octane, C8H18, is a substance in premium gasoline. How many moles are in 95.0 grams of octane?

2. What is the mass of 1.2 X 10^24 molecules of O2

3. How many moles are in 9.52g of MgCl2

4. How many moles are present in 180 grams of NaOH?

5. What is the mass in grams of 10.0 X 10^6 SO2 molecules?
1
Expert's answer
2019-11-11T07:48:07-0500

(1) 114.23 g of Octane = 1 mole

95 g. of Octane = (1/114.23) * 95 = 0.83 moles


(2) 6.023 * 1023 molecules = 32 g of O2

1.12 * 1024 molecules = (32/6.023*1023) * 1.12 * 1024 = 59.5 g.


(3) 95.2 g. of MgCl2 = 1 mole

9.52 g of MgCl2 = 0.1 mole


(4) 40 g of NaOH = 1 mole

180 g of NaOH = 180/40 = 4.5 moles


(5) 6.023 * 1023 molecules = 64 g

107 molecules = (64/6.022*1023 )* 107

= 10.62 * 10-16 g.